40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Jun 3 22:48:59 2018, johannes, Update, PSL, aux laser replacement 
    Reply  Mon Jun 4 02:52:52 2018, johannes, Update, PSL, aux laser replacement 
    Reply  Mon Jun 4 11:00:37 2018, gautam, Update, PSL, aux laser replacement 
       Reply  Tue Jun 5 02:06:59 2018, gautam, Update, PSL, aux laser first (NULL) results Close-But-no.jpg
Message ID: 13916     Entry time: Tue Jun 5 02:06:59 2018     In reply to: 13913
Author: gautam 
Type: Update 
Category: PSL 
Subject: aux laser first (NULL) results 

[johannes, gautam]

  1. Johannes aligned the single bounce off the ITM into the AUX fiber on the AS table, and also the AUX beam into the fiber on the PSL table.
    • Mode matching isn't spectaular anywhere in this chain.
    • But we have 2.6mW of light going into the SRM with the AOM deflection into the 1st order beam (which is what we send into the IFO) maximum.
  2. We set up some remote capabilities for the PLL and Marconi frequency (=PLL setpoint) control.
  3. Motivation was to try and lock DRMI, and look for some resonance of the AUX beam in the SRC.
    • We soon realized this was a way too lofty goal.
    • So we decided to try the simpler system of PRMI locked on carrier.
    • We were successfully able to sweep the Marconi setpoint in up to 20kHz steps (although we can only move the setpoint in one direction, not sure I know why now).
    • Then we decided to look for resonances of the AUX beam in the arm cavity.
    • Still no cigar broken heart
  4. Plus points:
    • PLL can be reliably locked remotely.
    • Marconi freq. can be swept deterministically remotely.
  5. Tomorrow:
    • Fix polarization issues. There is some low freq drift (~5min period) of the power incident on the fiber on the PSL table which we don't understand.
    • Verify MM into IFO and also into fiber at PSL table.
    • Do mode spectroscopy.

I was wondering why the PMC modulation sidebands are showing up on the control room analyzer with ~6dB difference in amplitude. Then I realized that it is reasonable for the cabling to have 6dB higher loss at 80 MHz compared to 20 MHz.

ELOG V3.1.3-